OpenStudy (anonymous):

Greg has 4 shirts: a white one, a black one, a red one, and a blue one. He also has two pairs of pants, one blue and one tan. What is the probability, if Greg gets dressed in the dark, that he winds up wearing the white shirt and tan pants? Show your work.

OpenStudy (anonymous):

c'mon u can do it

OpenStudy (doggy):

most of these answers can be found on google lol

OpenStudy (doggy):

like, super fast too XD

OpenStudy (anonymous):

u can't find google in exam

OpenStudy (doggy):

here's the exact same question, answered :P

OpenStudy (anonymous):

The table below shows the data for a random sample of fish that were collected from and later released into a lake: Type of Fish Catfish Grass Carp Bluegill Number of Fish 52 61 87 How many bluegills are estimated to be present if a sample of 800 fish is collected from the lake? 148 174 244 348

OpenStudy (kittiwitti1):

Catfish: 52 Grass Carp: 61 Bluegill: 87 Total amount of fish is 52+61+87=200 in the original count So, 87/200 are Bluegills. ?/800 in the new count are Bluegills. Proportion and simplify:\[\frac{87}{200}=\frac{x}{800}\rightarrow200x=87\times800\]Can you do it from here?

OpenStudy (anonymous):

is it d?

OpenStudy (kittiwitti1):

Yes. :)

OpenStudy (anonymous):

ok i have more questions :) next!

OpenStudy (anonymous):

??

OpenStudy (kittiwitti1):

Try a new post please.

OpenStudy (anonymous):

o k:)